OpenStudy (anonymous):

Verify the identity. (1 − sin^2 x)(tan^2 x + 1) = 1 Use the Pythagorean and Reciprocal Identities, and simplify. (1-sin^2 x)(tan^2 x1) = (?)(sec^2 x) = (?)(1/cos^2 x) = (?)

myininaya (myininaya):

Recall \[1-\sin^2(x)=\cos^2(x)\]

OpenStudy (anonymous):

(1-sin^x) = cos^2x tan^2x+1 = 1/cos^x Product of those two = 1.

OpenStudy (anonymous):

tan^2x+1 = 1/cos^2x I meant.